OpenStudy (anonymous):

solve 2x^2 + 7x + 3=0 solve using the quadratic formula

OpenStudy (anonymous):

simple:

OpenStudy (anonymous):

resolvent formula

OpenStudy (anonymous):

x=-0.5 x=-3

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

know resolvent form?

OpenStudy (anonymous):

i just attached the answer. :-)

OpenStudy (anonymous):

Using the quadratic formula, \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[x=\frac{-7 \pm \sqrt{7^2-4(2)(3)}}{2(2)}\] \[x=\frac{-7 \pm \sqrt{25}}{4}=\frac{-7 \pm 5}{4}=-\frac{1}{2}, -3\]

OpenStudy (anonymous):

You can also factor this using the grouping method: 2x^2+7x+3=0 2x^2+6x+x+3=0 2x(x+3)+(x+3)=0 (x+3)(2x+1)=0 x=-3, -1/2